# A Sparse Tensor Generator with Efficient Feature Extraction

### v2

@@ 6.1 Related Work
The previous works (24, 15) have considered the features for only one mode, assuming that the tensor is already sorted along that mode. However, real-world tensor sizes diverge significantly so that sizes in some modes reach millions while some are only orders of ten or even less. Therefore, considering the global values might result in losing some important information about the structure of the tensor.
Although several sparse matrix and graph generators are proposed in the literature (25, 26, 27, 28, 29), the studies on sparse tensor generators are very limited. Chi and Kolda (30) aim to produce low rank tensors; and Smith and Karypis (31) used their method to generate synthetic tensors whose values follow a Poisson distribution. Baskaran et al. (32) used MATLAB Tensor Toolkit (33) to generate synthetic sparse tensors but these are rather small tensors with less than one million nonzeros. Due to the deficiency of publicly available sparse tensors, Sun et al. (24) produced synthetic tensors by combining sparse matrices in Suite Sparse (34) collection such that the elements of matrices form the modes of tensors. However, real-world tensors are much sparser than real matrices and thus the structure of tensors generated by their method may differ significantly from the real ones.
@@ 6.2 Conclusion
In this study, we introduce two tools, FeaTensor and GenTensor, which we designed and developed to advance research in sparse tensor operations. FeaTensor is a feature extraction framework for sparse tensors that provides four different methods, prioritizing efficiency in extracting tensor features. This contribution is particularly valuable, as feature extraction itself is challenging and computationally expensive due to the large number of fibers and slices in real sparse tensors. We evaluate the performance of various feature extraction methods and observe that the methods introduced in this work outperform conventional approaches.
GenTensor is a smart sparse tensor generator that adheres to a comprehensive set of tensor features. Experimental results validate its effectiveness in mimicking real tensors, both in terms of sparsity patterns and tensor decomposition performance. A key advantage of GenTensor is its use of size-independent features, enabling the generation of tensors at different scales while preserving essential properties of real tensors. This capability facilitates the creation of large synthetic sparse tensor datasets that exhibit characteristics and behavior similar to real-world data. This is particularly useful for performance and scaling experiments involving tensor decomposition kernels that depend on tensor sparsity patterns. While this work focuses on sparsity patterns, we recognize that incorporating nonzero values and tensor rank poses additional challenges. Addressing these aspects would further enhance the quality of these tools, making it a promising direction for future research.

### v3

@@ 6.1 Related Work
The previous works (24, 15) have considered the features for only one mode, assuming that the tensor is already sorted along that mode. However, real-world tensor sizes diverge significantly so that sizes in some modes reach millions while some are only orders of ten or even less. Therefore, considering the global values might result in losing some important information about the structure of the tensor. Moreover, the feature extraction implementation of SpTFS is not publicly available in its official repository (25), which includes only the learning scripts and trained models. In contrast, our FeaTensor framework provides comprehensive multi-mode feature extraction without requiring any preprocessing. As a publicly available and parallelized tool, FeaTensor enables broader applicability and reproducibility, making it a more flexible and practical choice for large-scale sparse tensor analysis. Other widely used libraries for tensor decomposition or sparse storage optimization, such as SPLATT (18), ParTI (19), and HiCOO (12), do not provide any documented standalone feature extraction module either. Thus, to the best of our knowledge, FeaTensor represents the first open-source framework offering parallel, multi-mode feature extraction for sparse tensors.
Although several sparse matrix and graph generators are proposed in the literature (26, 27, 28, 29, 30), the studies on sparse tensor generators are very limited. Chi and Kolda (31) aim to produce low rank tensors; and Smith and Karypis (32) used their method to generate synthetic tensors whose values follow a Poisson distribution. Baskaran et al. (33) used MATLAB Tensor Toolkit (34) to generate synthetic sparse tensors but these are rather small tensors with less than one million nonzeros. Due to the deficiency of publicly available sparse tensors, Sun et al. (24) produced synthetic tensors by combining sparse matrices in Suite Sparse (35) collection such that the elements of matrices form the modes of tensors. However real-world tensors are much sparser than real matrices and thus the structure of tensors generated by their method may differ significantly from the real ones.
6.2 Limitations
To the best of our knowledge, FeaTensor and GenTensor are the first open-source tools with feature-preserving tensor generation and efficient multi-method feature extraction for sparse tensors. Despite their demonstrated utility, some limitations remain that offer opportunities for future enhancement. While this work primarily focuses on replicating structural sparsity patterns, it does not explicitly incorporate constraints on nonzero values or tensor rank. These aspects may affect downstream tasks such as classification, regression, or clustering, where the semantics of nonzero entries are important.
Additionally, although both FeaTensor and GenTensor are parallelized via OpenMP for shared-memory systems, their applicability to distributed-memory environments remains unexplored. Extending these tools to scale across nodes would broaden their utility in HPC settings. Addressing these aspects would further enhance the quality of these tools, making it a promising direction for future research.
GenTensor is designed with the flexibility to incorporate any distribution. However, its current version utilizes normal and log-normal distributions to determine the nonzero layout. Furthermore, the positions of the nonzeros, as well as the indices of nonzero slices and fibers, are selected uniformly, and nonzero values are drawn from a uniform distribution. As part of future work, we plan to extend GenTensor by exploring alternative distribution models for both layout and value generation.
In this work, the effectiveness of GenTensor has mainly been demonstrated through matching structural features and replicating CPD performance. These evaluation metrics were selected to capture both the structural fidelity and practical applicability of the generated tensors. Nonetheless, we acknowledge that other potential indicators could offer complementary perspectives. For instance, tensor rank or spectral characteristics might provide insight into latent structure fidelity. Moreover, evaluating the generated tensors under alternative decomposition models, such as Tucker, could reveal behavioral differences not captured by CPD alone. Graph-based similarity metrics also present a potential direction. While such alternatives may broaden the scope of evaluation, they often depend on specific applications or assumptions, which our current generic generator intentionally avoids. Exploring these dimensions in a targeted context remains a promising direction for future work.
@@ 6.3 Conclusion
In this study, we introduce two tools, FeaTensor and GenTensor, which we designed and developed to advance research in sparse tensor operations. FeaTensor is a feature extraction framework for sparse tensors that provides four different methods, prioritizing efficiency in extracting tensor features. It serves as the first publicly available, multi-mode, and parallel feature extraction framework for sparse tensors. This contribution is particularly valuable, as feature extraction itself is challenging and computationally expensive due to the large number of fibers and slices in real sparse tensors. We evaluate the performance of various feature extraction methods and observe that the methods introduced in this work outperform conventional approaches.
GenTensor is a smart sparse tensor generator that adheres to a comprehensive set of tensor features. Experimental results validate its effectiveness in mimicking real tensors, both in terms of sparsity patterns and tensor decomposition performance. A key advantage of GenTensor is its use of size-independent features, enabling the generation of tensors at different scales while preserving essential properties of real tensors. This capability facilitates the creation of large synthetic sparse tensor datasets that exhibit characteristics and behavior similar to real-world data. This is particularly useful for performance and scaling experiments involving tensor decomposition kernels that depend on tensor sparsity patterns. Moreover, experimental results demonstrate that GenTensor generalizes well to higher-order tensors and is robust to typical variations in the input feature set. The generator can therefore be reliably used in realistic benchmarking and modeling scenarios where input characteristics may not be precisely known.